Tr3:

Tomas bought some fruit at the market. He bought twice as many melons as apples. He bought 12 more melons than oranges. He bought 11 apples. How many pieces of fruit did Tomas buy? A. 43 B. 46 C. 23

Tr3:

@sillybilly123 ur pfp scares me

sillybilly123:

Tomas bought some fruit **f** at the market. He bought twice as many melons **m** as apples **a**: so m = 2a. He bought 12 more melons **m** than oranges **o**: so m = o + 12. He bought 11 apples **a**: so a = 11. And: f = m + a + o How many pieces of fruit **f** did Tomas buy?

sillybilly123:

a = 11 m = 2a = 22 m = o + 12 so o = 10 and f = m + a + o

sillybilly123:

👻